the simple interest on a certain sum for 3 1/2 years at 7% p.a is rs.225 more that the simple interest on the same sum for 4 years at 5% p.a find the sum
Answer by mananth(12270)


let principal be x
n=4 years
r= 5% p.a
SI= x*4*5/100
SI= x/5
II deposit
Principal =x
r=7
n= 3 1/2 years = 7/2 years
Si= (x*7*7)/(2*100)
SI=49x/200
II deposit interest -I deposit interest =225
49x/200 -x/5 =225
LCD = 200
multiply by 200 to eliminate the denominator
49x-40x=225*200
9x=225*200
x=225*200/9
x= Rs 5,000 the principal
m.ananth@hotmail.ca